ALGEBRA 1
Write in standard form: y+6=9/10(x+20)

Answer by mananth(16946)

y+6=9/10(x+20)
multiply by 10
10(y+6)=9(x+20)
10y+60=9x+180
-9x+10y=180-60
-9x+10y=120
...
m.ananth@hotmail.ca